3x + 1 = -7y
6x + 7y = 0
Solve for x and y.

Answer by chibisan(131)

solve sitmultaneously
3x + 7y = -1 -(1)
6x + 7y = 0 -(2)
(1)-(2)
(3x + 7y)(6x + 7y) = -1-0
3x - 6x = -1
-3x = -1
x = 1/3
finding y , subst x=1/3 to eq(1)
3(1/3) + 1 = -7y
2 = -7y
y = -2/7
x = 1/3 , y = -2/7 (ans)